The controller of Hartis Corporation estimates the amount of materials handling overhead cost that should be allocated to the company's two products using the data that are given below:

Wall Mirrors Specialty Windows
Total expected units produced 8,800 1,850
Total expected material moves 880 1,750
Expected direct labor-hours per unit 14 12


The total materials handling cost for the year is expected to be $17,404.50.


If the materials handling cost is allocated on the basis of material moves, the total materials handling cost allocated to the specialty windows is closest to: (Round your intermediate calculations to 2 decimal places.)

$5,573

$11,585

$8,902

$10,550
Answer #1

Answer

$11,585

Explanation:

The total materials handling cost = $17,404.50

Total expected material moves = 880+1750= 2,630

The materials handling cost per material move = $17,404.50/2,630 = $ 6.62

The materials handling cost for Specialty Windows = $ 6.62 * 1750 = $ 11,585

In case of any doubt, please comment.